OpenStudy (anonymous):

Write a matrix to represent the system. ð‘¥âˆ’ð‘¦+ð‘§=0ð‘¥âˆ’2ð‘¦âˆ’ð‘§=52ð‘¥âˆ’ð‘¦+2ð‘§=8

OpenStudy (perl):

$$ \Large{ x-y + z = 0 \\x - 2y - z = 5 \\2x -y + 2z = 8 \\~\\\iff\\~\\ \begin{bmatrix} 1& -1 & 1 \\1 & -2 & -1 \\ 2 & -1 & 2 \end{bmatrix} \cdot \begin{bmatrix} x\\y\\ z\end{bmatrix} = \begin{bmatrix} 0\\5\\ 8\end{bmatrix} } $$

OpenStudy (anonymous):

thank you!